# Reweighted Spectral Partitioning Works: Bounds for Special Graph Classes

### v3

@@ 1.1.1 Spectral Partitioning
Most relevant to this work is the method of spectral partitioning, which is the algorithm induced by the so-called “Cheeger inequality”, which is a well-known result in Spectral Graph Theory. For an undirected graph $G$ with $n$ vertices, let its adjacency matrix (or weight matrix) be denoted $A$, and the diagonal matrix formed from its degrees (or sums of incident weights) be denoted $D$. The Laplacian matrix of $G$ is $L(G):=D-A$. We denote the eigenvalues of $L(G)$ as $0=\lambda_{1}(G)\leq\lambda_{2}(G)\leq\lambda_{3}(G)\leq\cdots\leq\lambda_{n}(G)$. The study of these eigenvalues, and their relationships to combinatorial quantities, is a primary focus of the field of Spectral Graph Theory. The second-smallest Laplacian eigenvalue in particular ($\lambda_{2}(G)$) was dubbed the algebraic connectivity of a graph by Fiedler [Fie73], and is sometimes called the Fiedler value of a graph. As the name would suggest, algebraic connectivity has a relationship to a combinatorial form of connectivity. Namely, edge expansion:
@@
It is known that both sides of this inequality are asymptotically tight [HLW06]. A key feature of this theorem’s proof is that it is algorithmically constructive — the eigenvector associated with $\lambda_{2}(G)$ can be converted to a binary vector defining a cut via a threshold test, obtaining a sparse edge cut, giving the upper bound on edge expansion. $\lambda_{2}(G)$ can be computed (approximately) alongside a corresponding eigenvector in near-linear ($\tilde{\operatorname{\mathcal{O}}}(m)$) time [KMP11]. This is the algorithm called spectral partitioning. Applying this successively ($\operatorname{\mathcal{O}}(n)$ times), it can also be used to obtain edge-separators in $\tilde{\operatorname{\mathcal{O}}}(nm)$ time [ST96, ST07]. This algorithm is simple and fast, and, in fact, has guaranteed results (i.e., upper bounds on the sizes of the resulting separators) for many graph classes. These guarantees follow from upper bounds on $\lambda_{2}(G)$ for the classes, which have been found for a large number of graph classes (see Table 2, and see Appendix A for the definition of each graph class). Somewhat surprisingly, these bounds were all found after spectral partitioning had already seen widespread use as a heuristic.
{TAB}
(5,3pt,3pt)[5pt]—l—c—c——c—c—c—c—c—c—c—c— Graph class $\lambda_{2}(G)\lesssim$ Refs.
Planar $\dfrac{\Delta}{n}$ [ST96, ST07]
[…]
Triangulated [^1] Due to a subtle issue, this bound does not apply to general $g$ graphs, which is what the references claim. We will amend this issue as part of our work. $g$ $\dfrac{\text{poly}(\Delta)\cdot g}{n}$ [Kel04, Kel06b, Kel06a]
$K_{h}$-minor-free $\dfrac{\Delta\,(h\log h\log\log h)^{2}}{n}$ [BLR10, CF25, DP25]
RIG over $K_{h}$-minor-free $\dfrac{\Delta^{2}\,h^{6}\log h}{n}$ [Lee17]
$k$-ply neighbourhood system in $\mathbb{R}^{d}$ $\Delta\left(\frac{k}{n}\right)^{\frac{2}{d}}$ [ST96, ST07]
$k$-nearest-neighbour graph in $\mathbb{R}^{d}$ $\Delta\left(\frac{k}{n}\right)^{\frac{2}{d}}$ [ST96, ST07]
@@
Table 2: Best-known asymptotic upper bounds on $\lambda_{2}(G)$ for several graph classes. $n$ is always the number of vertices, and $\Delta$ is always the maximum degree. The notation $a\lesssim b$ means $a=\operatorname{\mathcal{O}}(b)$.
A downside of spectral partitioning is its dependence on the maximum degree $\Delta$ of the given graph. For instance, the resulting edge-separator size in planar graphs has a sub-optimal dependence on $\Delta$ compared to the best-known: it results in an edge-separator of size $\operatorname{\mathcal{O}}(\Delta\sqrt{n})$, but it is known that edge-separators exist of size $\operatorname{\mathcal{O}}(\sqrt{\Delta n})$ [Mil84, DDSV93]. In fact, this weaker bound is identical to what’s implied by the planar separator theorem – one can take the vertex separator of size $\operatorname{\mathcal{O}}(\sqrt{n})$ given by the planar separator theorem, and form an edge separator of size $\operatorname{\mathcal{O}}(\Delta\sqrt{n})$ from the edges incident to each vertex. This (very weakly) suggests that the techniques underpinning the eigenvalue bound could be more directly related to vertex-separators than edge-separators, in a more direct manner than what’s currently known.
1.1.2 Reweighted Spectral Partitioning
Since a larger value of $\lambda_{2}$ is associated with higher connectivity (and some other properties), one interesting question that could be asked is: Given a graph $G$, what is the weight distribution over the edges (under some normalization constraints) that maximizes $\lambda_{2}$? Intuitively, in some cases, this might let us reweigh the edges of a graph to improve its connectivity. This question was first posed by Boyd, Diaconis and Xiao, who recognized it as a variant of the fastest-mixing Markov chain problem [BDX04]. They refer to the extremal value as $\lambda_{2}^{*}(G,\pi)$ for an undirected connected graph $G$, and a probability mass function $\pi:V\to[0,1]$ of the vertices, and it is formulated as the following optimization problem (also called the maximum reweighted spectral gap):
$\begin{array}[]{r c c r c l l}\lambda_{2}^{*}(G,\pi)&:=&\max_{P\geq 0}&\lambda_{2}(I-P)\\ &&\textup{subject to}&P(u,v)&=&0&\forall uv\not\in E\cup\{vv:v\in V(G)\}\\ &&&\sum_{v\in V}P(u,v)&=&1&\forall u\in V\\ &&&\pi(u)P(u,v)&=&\pi(v)P(v,u)&\forall uv\in E\cup\{vv:v\in V(G)\}\end{array}$
@@ 1.1.2 Reweighted Spectral Partitioning
$P$ is a reweighted adjacency matrix of $G$ (with self-loops added) normalized to represent a reversible Markov chain with stationary distribution $\pi$. Hence, $I$ is the weighted degree-matrix for the weighted graph, $I-P$ is the Laplacian matrix of the weighted graph, and $\lambda_{2}(I-P)$ is its second-smallest eigenvalue. We will restrict ourselves to discussing the case where $\pi$ is the uniform distribution, which we denote simply as $\lambda_{2}^{*}(G)$. Before moving on, we make an important observation:
Observation 1.2.
For a graph $G$ with maximum degree $\Delta$, $\lambda_{2}(G)\leq\Delta\cdot\lambda_{2}^{*}(G)$.
Boyd, Diaconis, and Xiao also realize their optimization problem as a Semidefinite Program meeting Slater’s condition (therefore allowing it to be approximately solved in polynomial time via e.g. [ZL18]), and in the process they derive a dual formulation (better stated and properly extracted in follow-up work by Roch [Roc05]):
$\begin{array}[]{r c c r c l l}\gamma^{(d)}(G)&:=&\min_{\begin{subarray}{c}f:V\to\mathbb{R}^{d}\\ y:V\to\mathbb{R}_{\geq 0}\end{subarray}}&\sum_{v\in V}y(v)\\ &&\textup{subject to}&\sum_{v\in V}f(v)&=&\hskip 10.00002pt\overline{0}\\ &&&y(u)+y(v)&\geq&\hskip 10.00002pt\frac{||f(u)-f(v)||_{2}^{2}}{\sum_{x\in V}||f(x)||_{2}^{2}}&\forall uv\in E\end{array}$ (1)
A strong duality proof gives that $\lambda_{2}^{*}(G)=\gamma^{(n)}(G)$. Note that we have intentionally added the choice of dimension $d$ (rather than fixing $d=n$) for later use. Other methods of computing $\lambda_{2}^{*}(G)$ include a particularly simple and practical sub-gradient method [BDX04] and a matrix-multiplicative weight-update method that runs in near-linear time if $\lambda_{2}^{*}$ is very large and weighted in a particular way [LTW24, Section 5].
One useful bound on $\lambda_{2}^{*}(G)=\gamma^{(n)}\leq\gamma^{(1)}(G)$ is:
Observation 1.3.
For any graph $G$, $\gamma^{(1)}(G)\lesssim 1$.
Proof.
If $n$ is even, assign $y$ to be $\frac{2}{n}$ for every vertex, and assign $f$ to be 1-1- 1 for half the vertices and 111 for the other half of the vertices, so then $\gamma^{(1)}(G)\leq 2\lesssim 1$. If $n$ is odd (and at least 333), assign $y$ to be $\frac{2}{n-1}$ for every vertex, and assign $f$ to be an even split of 1-1- 1 and 111 for all but one vertex, which is assigned 0, resulting in $\gamma^{(1)}(G)\leq 2+\frac{2}{n-1}\lesssim 1$. ∎
A series of work has recently resulted in a Cheeger-style inequality relating vertex expansion with $\lambda_{2}^{*}$:
@@
For an undirected graph $G$ with maximum degree $\Delta$, and any probability distribution $\pi$ on $V$,
$\frac{\psi(G)^{2}}{\log\Delta}\lesssim\lambda_{2}^{*}(G)\lesssim\psi(G).$
This is the single most important known theorem involving $\lambda_{2}^{*}$. Like the proof of Cheeger’s inequality, Theorem 1.4 is also algorithmically constructive, and the implied randomized polynomial-time algorithm that converts a solution to $\lambda_{2}^{*}(G)$ (or more specifically, $\gamma^{(n)}(G)$) into a bounded-ratio vertex-cut upper-bounding $\psi(G)$ is the family of methods we call reweighted spectral partitioning. We call this a “family” of methods since a few different methods for each step of the proof (with slightly different constants) have been given. Moreover, we will provide yet another construction for one step with slightly different bounds later in this work to arrive at a “refined” Cheeger inequality for vertex expansion in Section 5.
The “easy” direction Theorem 1.4 was shown by Roch [Roc05]. A weaker form of the hard direction was given by Olesker-Taylor and Zanetti [OTZ22], which was quickly refined by Jain, Pham, and Vuong [JPV22], and independently by Kwok, Lau, and Tung [KLT22] (who also handled the weighted case, which we omit). Our outline will follow the techniques of Kwok, Lau, and Tung. There are two separate results that together imply both sides of the inequality:
Theorem 1.5 ([OTZ22, Theorem 2.10]).
For a graph $G$,
$\psi(G)^{2}\lesssim\gamma^{(1)}(G)\lesssim\psi(G).$
This result is generalized by Kwok, Lau, and Tung for the weighted case [KLT22, Theorem 3.15]. Their proof is notable for our case of uniform weights too, since it implies a very simple and practical algorithm for this step.
Theorem 1.6 ([KLT22, Proposition 3.14]).

### v4

@@ 2.1 Fiedler Bounds and the Cheeger Inequality
For an undirected graph $G$ with $n$ vertices, let its adjacency matrix (or weight matrix) be denoted $A$, and the diagonal matrix formed from its degrees (or sums of incident weights) be denoted $D$. The Laplacian matrix of $G$ is L(G):=DAL(G):=D-A, and its eigenvalues are denoted $0=\lambda_{1}(G)\leq\lambda_{2}(G)\leq\lambda_{3}(G)\leq\cdots\leq\lambda_{n}(G)$. The second-smallest Laplacian eigenvalue ($\lambda_{2}(G)$) was named the algebraic connectivity of a graph by Fiedler [Fie73], and is sometimes called the Fiedler value of a graph. As the name would suggest, algebraic connectivity is related to a combinatorial definition of connectivity:
For a graph $G=(V,E)$, the edge expansion of a set $S\subset V$ with $|S|\leq\frac{n}{2}$ is defined as $\phi(S):=\frac{|\{uv\in E\cap S\times(V\setminus S)\}|}{|S|}$, and the edge expansion of the graph $G$ is defined as $\phi(G):=\min_{|S|\leq\frac{n}{2}}\phi(S)$.
The edge expansion $\phi(G)$ and the Fiedler value $\lambda_{2}(G)$ are related by the following important result:
@@
This inequality is very similar to Theorem 1.4 in a number of ways. In particular, there is a simple near-linear-time algorithm, called spectral partitioning, that can produce a set $S$ with $\phi(S)\lesssim\sqrt{\Delta\lambda_{2}(G)}$ [KMP11]. Very similarly to our results, one can obtain direct bounds on $\lambda_{2}(G)$ for several special graph classes (see Table 2). Consequently, this algorithm can be used to produce good cuts, and even near-optimal balanced vertex-separators if the maximum degree is constant.
@@
Table 2: Best-known asymptotic upper bounds on $\lambda_{2}(G)$ for several graph classes, for $n$-vertex graphs of maximum degree $\Delta$.
2.2 The Fastest-Mixing Markov Chain Problem and the Cheeger-Style Inequality for Vertex Expansion
Since the Fiedler value $\lambda_{2}(G)$ is associated with higher connectivity, it is natural to ask if different weights along the edges of the graph can result in higher values of $\lambda_{2}(G)$. Without any constraints on the weights, the value of $\lambda_{2}(G)$ is unbounded. A natural normalization constraint is to require that the total weight around each vertex is at most $1$. Boyd, Diaconis and Xiao studied exactly this problem, which they called the fastest-mixing (reversible) Markov chain problem [BDX04]:
$\begin{array}[]{r c c r c l l}\lambda_{2}^{*}(G)&:=&\max_{P\geq 0}&\lambda_{2}(I-P)\\ &&\textup{subject to}&P(u,v)&=&0&\forall uv\not\in E\cup\{vv:v\in V(G)\}\\ &&&\sum_{v\in V}P(u,v)&=&1&\forall u\in V\\ &&&P(u,v)&=&P(v,u)&\forall uv\in E\cup\{vv:v\in V(G)\}\end{array}$
@@ 2.2 The Fastest-Mixing Markov Chain Problem and the Cheeger-Style Inequality for Vertex Expansion
Here $P$ is a reweighted adjacency matrix of $G$ (with self-loops) constrained to produce a reversible Markov chain with a uniform stationary distribution. Hence, $I$ is the degree-matrix for the weighted graph, and IPI-P is the Laplacian matrix, so $\lambda_{2}(I-P)$ is the Fiedler value.
33footnotetext: Due to some subtle issues, this bound does not apply to general genus-$g$ graphs, which is what the references claim. We both discuss and amend this issue in Section 5.3.
There are two simple and important relationships relating $\lambda_{2}(G)$, $\lambda_{2}^{*}(G)$, and $\gamma^{(n)}(G)$: First, it can be shown that $\lambda_{2}^{*}(G)=\gamma^{(n)}(G)$ by a duality argument [BDX04, Roc05]. Second, $\lambda_{2}(G)\leq\Delta\cdot\lambda_{2}^{*}(G)$. Therefore, Theorem 1.5 actually generalizes the known bounds on $\lambda_{2}(G)$. In fact, one of the resulting bounds is new:
Corollary 2.2.
If $G$ is a genus-$g$ graph, then its Fiedler value satisfies $\lambda_{2}(G)\lesssim\frac{g\cdot\Delta\log\Delta}{n}$.
We will henceforth use $\gamma^{(n)}(G)$ instead of $\lambda_{2}^{*}(G)$ since they are equal.
There has been significant prior work towards relating $\gamma^{(n)}$ to vertex-expansion. In particular, our Theorem 1.4 is a refinement of the following result:
@@
For an undirected graph $G$ with maximum degree $\Delta$
$\frac{\psi(G)^{2}}{\log\Delta}\lesssim\gamma^{(n)}(G)\lesssim\psi(G).$
This line of work is what gives rise to the name reweighted spectral partitioning. In particular, the most recent improvements to this result were in the dimension-reduction step of the algorithm, improving a $\log n$ factor to the current $\log\Delta$ factor [JPV22, KLT22]. Our result further improves this dimension-reduction step to get a $\min\{\alpha(G)^{2},\log\Delta\}$ factor.
2.3 Bounds via Geometry with Ball-Intersection Graphs
Our first family of methods for obtaining upper bounds makes use of ball-intersection graphs. These are a very natural object to consider: For any feasible solution to $\gamma^{(d)}(G)$, we will show that $f,y$ induce a ball-intersection graph $H=(V,F)$ with centre $f(v)$ and radius $2\sqrt{y(v)}$ for each vertex $v\in V$. In fact, we will reformulate $\gamma^{(d)}(G)$ into a new form that more directly relates to such a representation, at only a constant-factor loss.
We will use this form to give upper bounds on $\gamma^{(d)}(G)$ for planar graphs, genus-$g$ graphs, ball-intersection graphs of bounded ply, and $k$-nearest-neighbour graphs. Specifically, we will obtain the following bounds:
Proposition 2.4.